OpenStudy (anonymous):

How many ways are there to choose eight coins from a piggy bank containing 100 identical pennies and 80 identical nickels

OpenStudy (kropot72):

8P give 1 way. 8N give 1 way. 1P 7N give: \[\frac{8!}{7!}=8\ ways.\] 2P 6N give: \[\frac{8!}{2!6!}=28\ ways.\] 3P 5N give: \[\frac{8!}{3!5!}=56\ ways.\] 4P 4N give: \[\frac{8!}{4!4!}=70\ ways.\] 5P 3N give 56 ways. 6P 2N give 28 ways. 7P 1N give 8 ways. The total number of of ways to choose eight coins is give by: 1 + 1 + 8 +28 +56 + 70 + 56 +28 + 8 = you can calculate.